1 ) If tanA=a/b , then (asinA+bcosA)/(asinA-bcosA)=?

Grade:12

2 Answers

sanatan sharma
12 years ago

if i divide the numerator and denominator by cosA we will get 

= ((asinA/cosA)+b)/((asinA/cosA)-bcosA)

= (atanA + b)/(atanA - b)

= (a2 + b2)/(a- b2)

Muntasir Mahi
4 years ago
tanA=a/b
=sinA/cosA=a/b
=a/b×sinA/cosA=a/b×a/b
=asinA/bcosA=a2/b2
=asinA+bcosA/asina-bcosA=a2+b2/a2-b2